if x=(2+root3) find the value of xsquare+1/x square

Grade:9

1 Answers

Subhajit Gorai
3 years ago
x=2+3^(1/2)
Find x^2+1/x^2?
 
I= x^2+1/x^2
I+2= x^2+1/x^2+2    [ADDING 2 ON BOTH SIDES]
I+2={x+1/x}^2
I+2={(x^2+1)/x}^2
PUTTING THE VALUE OF x=2+3^(1/2):
I+2=4^2
I=14(ans)